### v2

# Mixed updating in structured populations

Abstract
Evolutionary graph theory (EGT) studies the effect of population structure on evolutionary dynamics. The vertices of the graph represent the $N$ individuals. The edges denote interactions for competitive replacement. Two standard update rules are death-Birth (dB) and Birth-death (Bd). Under dB, an individual is chosen uniformly at random to die, and its neighbors compete to fill the vacancy proportional to their fitness. Under Bd, an individual is chosen for reproduction proportional to fitness, and its offspring replaces a randomly chosen neighbor. Here we study mixed updating between those two scenarios. In each time step, with probability $\delta$ the update is dB and with remaining probability it is Bd. We study fixation probabilities and times as functions of $\delta$ under neutral evolution and constant selection. Despite the fact that fixation probabilities and times can be increasing, decreasing, or non-monotonic in $\delta$, we prove nearly all unweighted undirected graphs have short fixation times and provide an efficient algorithm to estimate their fixation probabilities. Finally, we prove exact formulas for fixation probabilities on cycles, stars, and more complex structures and classify their sensitivities to $\delta$.
00footnotetext: Corresponding author: David A. Brewster (dbrewster@g.harvard.edu)
1 Introduction
In evolutionary graph theory, population structure is described by a weighted or unweighted graph together with an update rule. Individuals occupy the vertices of the graph, and edges specify potential replacement interactions. The neighbors of a vertex indicate the possible locations where offspring can be placed. If the graph is weighted, some replacement events occur with higher probability than others; a self-loop indicates that an offspring may replace its parent. Reproduction is typically assumed to be asexual. A well-mixed population is represented by an unweighted complete graph with self-loops.
Evolutionary graph theory can be used to study neutral evolution [1, 2, 3], constant selection [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or frequency dependent selection [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The latter is an approach for evolutionary game theory in structured populations. In this paper, we focus on neutral evolution and constant selection.
Two update rules that have been widely considered are death-Birth (dB) updating and Birth-death (Bd) updating. For death-Birth (dB) updating, an individual is selected for death uniformly at random among the population. Then the neighbors compete for the empty site proportional to their fitness and the weight of the link that leads from them to the empty site. For Birth-death (Bd) updating, an individual is selected for reproduction proportional to its fitness; then the offspring replaces a neighbor with a probability that is proportional to the weight of the edge that leads to this neighbor (see [32]).
Past work in evolutionary graph theory has yielded a rich set of results regarding fixation probabilities and fixation times [33, 34, 6, 7, 8, 9, 10, 12, 35, 36, 37]. The fixation probability is the probability that a newly introduced mutant takes over the entire population. The fixation probability can be calculated for a mutant that starts in a specific place or averaged over all initial placements. The unconditional fixation time (also known as the absorption time) refers to the average time it takes for a population to reach homogeneity after the introduction of a mutant. The (conditional) fixation time is the average time it takes for the new mutant to reach fixation.
There are population structures that amplify or suppress selection [6, 7, 8, 9, 10, 35, 36, 38]. An amplifier increases the fixation probability of an advantageous mutant and reduces the fixation probability of a deleterious mutant. A suppressor does the opposite. There are results on fast and strong amplifiers [39, 13]. There are population structures that can maintain diversity for a long time or eliminate diversity quickly [3, 40]. In the case of a well-mixed population, the classic Moran process is recovered, regardless of whether Bd or dB updating is used [41]. However, it is typically the case that Bd and dB updating leads to different results in structured populations [42, 43, 44].
In this paper, we study $\delta$-mixed updating: in any one time step, we use death-Birth (dB) with probability $\delta$ and we use Birth-death (Bd) updating with probability $1-\delta$. There have been attempts to bridge the two worlds of dB and Bd, though those works have limited exploration into the enormous universe of the possible population structures and do not analyze fixation times [26, 13, 45]. In [26], the authors study $\delta$-mixed updating and prove that for the Prisoners’ Dilemma on a $k$-regular graph, cooperation is favored when $b/c>k/\delta$ where $b$ is the benefit from a cooperator, $c$ is the cost a cooperator pays, and $b>c>0$; if $\delta=0$, cooperation is never favored. Essentially, there is a smooth interpolation for the result of [46] between dB updating and Bd updating. In [13], the authors study graphs that maintain amplification as $\delta$ moves away from 0; they find that unlike the case in pure Bd (i.e., $\delta=0$), when $\delta>0$, amplification can only be a transient phenomenon in regards to the selection advantage. In [45], the authors use a continuous-time updating dynamics such that when normalized, there is a natural death rate $\Delta$ that triggers a death-Birth event, and there is a birth rate $\varphi$ that triggers a Birth-death event where $\varphi$ is the average fitness of the current population. Then their continuous-time model is $\delta$-mixed updating with $\delta=\Delta/(\varphi+\Delta)$. The authors find that stars on $N$ vertices with self-loops have a critical threshold at $\Delta\sim 1/\sqrt{N}$: stars are amplifiers of selection under uniform initialization when $\delta\ll 1/\sqrt{N}$ and suppressors of selection when $\delta\gg 1/\sqrt{N}$. We study fixation probability and fixation time on unweighted undirected graphs as a function of $\delta$ in the case of neutral evolution and constant selection.
2 Model
Figure 1: Mixed $\delta$-updating on a graph. Initially, NN individuals are present in the population with one mutant. At each time step, a death-Birth step occurs with probability $\delta$. With remaining probability, a Birth-death step occurs. The thick arrows on the edges of the graph are oriented such that the arrowheads indicate the possible location(s) of death; the tails indicate the possible location(s) of birth for each time step.
2.1 Setup
Let $G=(V,E)$ be a connected undirected graph on $N:=|V|$ vertices. The number of neighbors of a vertex in a graph is called its degree. The degree of vertex $u\in V$ is denoted $\deg(u)$. On each vertex of $G$ lies an individual that is either a wild type or a mutant type. Mutants have relative fitness $r\geq 1$ and wild-types have fitness $1$. When $r=1$, the evolution is neutral; when $r>1$, mutants have a selective advantage (constant selection).
2.2 Dynamics
Initially, only a subset $S_{0}\subseteq V$ of the graph is occupied by mutants. Then at discrete time steps $t=1,2,\ldots$, a coupled birth & death occurs. We call subset of mutants $S_{t}\subseteq V$ the configuration of the process after $t$ steps. The ordering of these coupled events matters. There are two possible orderings we consider.
death-Birth (dB). A death-Birth (dB) step proceeds as follows. We first select an individual $v\in V$ for death uniformly at random. Then we select an incoming neighboring vertex $u\in V$ of $v$ to give birth proportional to fitness. Then the individual at location $u$ places a copy of itself at the recent vacancy of location $v$.
Birth-death (Bd). A Birth-death (Bd) step proceeds as follows. We first select an individual $u\in V$ for birth proportional to fitness. Then we select an outgoing neighboring vertex $v\in V$ of $u$ to die uniformly at random. The individual at location $u$ then places a copy of itself (and thus of its type) at the recent vacancy of location $v$.
We will refer to dB and Bd as pure updating rules since the updating rule is fixed for each step. Next, we define an interpolation of the pure updating rules.
Mixed $\delta$-updating. Let $\delta\in[0,1]$ be the probability a given time step will be a dB step. With remaining probability, the step will be a Bd step. We only consider the case when the value of $\delta$ is constant throughout the entire process. Note that we recover pure dB updating when $\delta=1$ and recover pure Bd updating when $\delta=0$. When $\delta<1/2$, the process is Bd-biased; when $\delta>1/2$, the process is dB-biased. When $\delta=1/2$, we refer to the process as unbiased. See Figure 1 for an illustration of each step of the process.
2.3 Fixation probability and time
Eventually, the population becomes entirely mutant—we call this fixation—or entirely wild-type—we call this extinction. In either case, we say that absorption has occurred. We denote the probability of fixation given mutant fitness $r\geq 1$ and initial set $S_{0}\subseteq V$ as $\mathsf{fp}_{r}^{\delta}(G,S_{0})$. We also denote the uniform initialization fixation probability as $\mathsf{fp}_{r}^{\delta}(G):=\frac{1}{N}\sum_{u\in V}\mathsf{fp}_{r}^{\delta}(G,\{u\})$. When dealing with times, we focus solely on the expected time rather than the entire distribution. For an initial mutant set $S_{0}\subseteq V$, we denote the (expected) time until fixation, extinction, and absorption as $\mathsf{FT}_{r}^{\delta}(G,S_{0})$, $\mathsf{ET}_{r}^{\delta}(G,S_{0})$, and $\mathsf{AT}_{r}^{\delta}(G,S_{0})$, respectively. Since we are interested in upper bounds for times, we drop the mutant initial set from the notation for time to denote the maximum time over all possible initial mutant sets. Thus, for example, $\mathsf{FT}_{r}^{\delta}(G):=\max_{S_{0}\subseteq V}\mathsf{FT}_{r}^{\delta}(G,S_{0})$. In the notations for probabilities and times, we use $\delta=\textrm{dB}$ to mean $\delta=1$ and $\delta=\textrm{Bd}$ to mean $\delta=0$. When the fitness is neutral ($r=1$), we write $\mathsf{fp}_{r=1}^{\delta}$, $\mathsf{FT}_{r=1}^{\delta}$, etc.
3 Results
Figure 2: Fixation probabilities, $\rho$, of various graphs under mixed $\delta$-updating for $N=7$ in neutral evolution. In both figures, the dashed line at $1/N$ is the fixation probability of the Moran process (well-mixed) in neutral evolution. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. a, i) A star graph with the initial mutant location at the center. Relative to $\rho$ under the unbiased process, Bd-biased processes have a lower $\rho$ but dB-biased processes have a higher $\rho$. a, ii) A star graph with the initial mutant location at a leaf. Relative to $\rho$ under the unbiased process, Bd-biased processes have a higher $\rho$ but dB-biased processes have a lower $\rho$. a, iii) A graph composed of a star and two additional edges. Relative to $\rho$ under the unbiased process, both pure Bd and pure dB processes have a lower $\rho$. However, panel b, and iv) shows that the graph from a, iii) has higher $\rho$ than the unbiased process when $\delta$ is slightly smaller than $1/2$. This example shows that $\rho$ can be non-monotonic in $\delta$. Further, it is not possible that both pure Bd and pure dB have higher $\rho$ than the unbiased process (see Lemma D.1).
3.1 Neutral evolution
Suppose that $G=(V,E)$ is graph and the initial set of mutants reside at $S_{0}\subseteq V$. For the pure Bd process in neutral evolution, it is known that $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,S_{0})\propto\sum_{u\in S_{0}}1/\deg(u)$ [47]; for the pure dB process, it is known that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,S_{0})\propto\sum_{u\in S_{0}}\deg(u)$ [44]. Due to the additivity of the fixation probability in the pure processes, $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G)=\mathsf{fp}_{r=1}^{\mathrm{dB}}(G)=1/N$.
First, we prove that under general $\delta$, the fixation probability is similarly additive in the initial mutant set like it is in the pure Bd and dB processes; that is $\mathsf{fp}_{r=1}^{\delta}(G,S_{0})=\sum_{u\in S_{0}}\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ (see Lemma B.1). Since $\mathsf{fp}_{r=1}^{\delta}(G)=(1/N)\cdot\sum_{u\in V}\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$, additivity easily yields the following result.
Theorem 3.1 (Uniform initialization).
Averaging uniformly over all initial conditions, the fixation probability of a neutral mutant is $1/N$ on any graph for any value of delta.
In our notation, Theorem 3.1 means that $\mathsf{fp}_{r=1}^{\delta}(G)=1/N$ for any graph $G$ and any value of $\delta$. Thus mixed $\delta$-updating under uniform initialization is well-behaved. Let $u\in V$. It is known that if one of $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})$ or $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\})$ is larger than $1/N$, then the other is at most $1/N$ (see Lemma D.1). Surprisingly, we prove that no matter the fixation probabilities of the pure processes with initial mutant location $u$, the unbiased process always has fixation probability $1/N$ when the initial mutant location is $u$. This means that even if both pure processes have fixation probability less than $1/N$ for initial mutant location $u$, the unbiased process has a higher fixation probability than both the pure processes for initial mutant location $u$.
Theorem 3.2 (Arbitrary initialization).
If $\delta=1/2$ (which is the unbiased case) then the fixation probability of a neutral mutant starting in an arbitrary location is $1/N$ on any graph.
In our notation, Theorem 3.2 states that for $G=(V,E)$ and $u\in V$ we have $\mathsf{fp}_{r=1}^{\delta=1/2}(G,\{u\})=1/N$. As demonstrated by Figure 2, the quantity $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ as a function of $\delta$ can be increasing, decreasing, concave, convex, or non-monotone. Despite this possible strange behavior, we prove that the fixation probabilities on regular graphs are not affected by $\delta$.
Theorem 3.3 (Fixation probability of regular graphs).
The fixation probability of a neutral mutant starting in an arbitrary location on a regular graph is $1/N$ for any value of $\delta$.
In fact, we prove something stronger: there is a broad class of graphs whose fixation probabilities are monotone in $\delta$. Such graphs—called bidegreed—have a vertex cut such that the vertices in the cut all have the same degree and the vertices outside the cut all have the same degree.
Definition 3.1 (Bidegreed graph).
A bidegreed graph is an undirected graph with the following property: for each vertex the number of neighbors is either $d_{1}$ or $d_{2}$.
If $d_{1}=d_{2}$ then the graph is regular. The following statements are proven in Appendix E.1.
Theorem 3.4 (Bidegreed graphs).
On a bidegreed graph with $d_{1}<d_{2}$, the fixation probability of a neutral mutant starting on a vertex with degree $d_{1}$ is a strictly decreasing function of $\delta$. The fixation probability of a neutral mutant starting on a vertex with degree $d_{2}$ is a strictly increasing function of $\delta$.
Let $n_{i}$ denote the number of vertices in a bidegreed graph $G$ with degree $d_{i}$ for $i=1,2$. We denote $\rho_{i}(\delta)$ as $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ where $u$ is a vertex with degree $d_{i}$. The sensitivity of a vertex with degree $d_{i}$ is $s_{i}(\delta):=|\rho^{\prime}_{i}(\delta)|$. In addition to Theorem 3.4, we prove that for $i=1,2$,
  1. 1.
    if $n_{1}>n_{2}$, $s_{i}$ has its maximum at $\delta=1$ and minimum at $\delta=0$;
  2. 2.
    if $n_{1}<n_{2}$, $s_{i}$ has its maximum at $\delta=0$ and minimum at $\delta=1$;
  3. 3.
    if $n_{1}=n_{2}$, then $s_{i}$ is constant with respect to $\delta$.
See Figure 3 for examples of bidegreed graphs and their behaviors with respect to $\delta$. A star graph is a canonical example of a bidegreed graph that has extreme degrees (see Figure 4).
Figure 3: Examples of $(d_{1},d_{2})$-bidegreed graph families (a-f) and their sensitivities in neutral evolution to initial mutant location at $\delta=0,1$ (g-j). Vertices with degree $d_{1}$ are represented as circles whereas vertices with degree $d_{2}\geq d_{1}$ are represented as squares. For the examples depicted, $d_{2}$ is strictly greater than $d_{1}$, but it is also possible for $d_{2}=d_{1}$ per Definition 3.1. Let $n_{1}$ denote the number of vertices with degree $d_{1}$ and $n_{2}$ denote the number of vertices with degree $d_{2}$. The value $\rho_{i}(\delta)$ is the derivative of $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ with respect to $\delta$ where $u$ is a vertex with degree $d_{i}$. a, a star graph, $S_{N}$; the $n_{2}=1$ center vertex has degree $d_{2}=6$ while the $n_{1}=6$ periphery vertices have degree $d_{1}=1$ for a total of $N=n_{1}+n_{2}=6+1=7$ vertices. b, a path graph, $P_{N}$; the intermediate vertices have degree 22 while the endpoint vertices have degree $1$. c, a bipartite-biregular graph; the vertices on the left side have degree $\ell=3$ while the vertices on the right side have degree $r=2$; there are $N=10$ vertices. d, a fan graph $F_{N}$; the center vertex has degree $n=2b=6$ while each of the $2$ vertices on the $b$ blades have degree $2$, for a total of $N=2b+1=7$ vertices. e, a wheel graph $W_{N}$; the center vertex has degree $n=8$ while the $n$ vertices on the spoke ends have degree $3$, for a total of $N=n+1=9$ vertices. f, a book graph $B_{N}$; each of the $p=3$ pages has 22 page corners with degree 22; the 22 seam endpoints have degree $p+1$, for a total of $N=2p+2=8$ vertices. g, A contour plot in the $(\eta,\kappa)$ space where $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ are dimensionless parameters. The lines with shapes denote the values of $\eta=n_{1}/n_{2}$ and $\kappa=d_{2}/d_{1}$ for the various bidegreed graphs in a-f. Note that the path graph in our examples since it has $n_{1}\leq n_{2}$ for $N\geq 4$; thus as NN increases, $\eta$ decreases. The intensity of the colorbar denotes the magnitude of $\rho^{\prime}_{1}(\delta=0)$, h, $\rho^{\prime}_{2}(\delta=0)$, i, $\rho^{\prime}_{1}(\delta=1)$, and j, $\rho^{\prime}_{2}(\delta=1)$.
Key to our proof of Theorem 3.4 is an explicit formula for the fixation probability on a bidegreed graph for any $\delta$. The proof of the following theorem is in Appendix E.
Theorem 3.5 (Bidegreed fixation probabilities).
On a bidegreed graph with $d_{1}\leq d_{2}$ and for all $\delta$, the fixation probability of a neutral mutant starting on a vertex with degree $d_{1}$ is proportional to $1$. The fixation probability of a neutral mutant starting on a vertex with degree $d_{2}$ is proportional to $\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}$.
A star graph $S_{N}$ is a canonical example of a bidegreed graph. With center $c$ and n=N1n=N-1 leaves $\ell_{1},\ldots,\ell_{n}$, Theorem 3.5 gives a formula for the fixation probability for any $\delta$. We can also state it explicitly.
Theorem 3.6 (Fixation probability on stars, neutral evolution).
For a star $S_{N}$ with center $c$ and n=N1n=N-1 leaves, for each $i\in\{1,\ldots,n\}$ and all $\delta\in[0,1]$,
$\mathsf{fp}_{r=1}^{\delta}(S_{N},\{\ell_{i}\})=\frac{1+(1-\delta)(N-2)}{(1-\delta)(N-2)^{2}+2(N-1)}\quad\text{and}\quad\mathsf{fp}_{r=1}^{\delta}(S_{N},\{c\})=\frac{1+\delta(N-2)}{(1-\delta)(N-2)^{2}+2(N-1)}.$ (1)
The proof is in Section 7 in the SI.
Figure 4: Fixation probabilities on a star graph with $N=10$ vertices in neutral evolution. The various curves represent the various values of $\delta$. The dashed lines are the theoretical fixation probabilities from Theorem 3.6. The translucent regions represent the $95\%$ confidence intervals for the average of $10^{4}$ simulations of the mixed $\delta$-updating process; a, Fixation probabilities as a function of $\delta$ when starting with a mutant in the center. b, Same as a) except a mutant is initially on a leaf.
Similar to fixation probabilities, fixation times in the mixed $\delta$-updating process can exhibit a large range of behaviors relative to $\delta$ (see Figure 5). For graph families $(G_{N})_{N\geq 1}$, it is known $\mathsf{AT}_{r=1}^{\textrm{Bd}}(G_{N})$ is $O(N^{6})$ [1]. For pure dB, it is known $\mathsf{AT}_{r=1}^{\textrm{dB}}(G_{N})$ is $O(N^{5})$ [48]. Relating fixation times to absorption times, the law of total expectation on the absorption time gives the bound $\mathsf{FT}_{r=1}^{\delta}(G,S_{0})\leq\mathsf{AT}_{r=1}^{\delta}(G,S_{0})/\mathsf{fp}_{r=1}^{\delta}(G,S_{0})$ for any $S_{0}\subseteq V$ and $\delta\in[0,1]$.
Figure 5: a, Fixation times, $T$, of various graphs under mixed $\delta$-updating for $N=7$ in neutral evolution. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. i) The fixation time increases and then decreases relative to $\delta$. ii) The fixation time decreases and then increases relative to $\delta$. iii) The fixation time always increases relative to $\delta$. iv) The fixation time always decreases relative to $\delta$.
We prove that the absorption and fixation times are well-behaved in the unbiased process.
Theorem 3.7 (Fixation time).
If $\delta=1/2$ (which is the unbiased case), then the fixation time of a neutral mutant starting in an arbitrary location on any graph of size $N$ is at most $N^{5}/4$.
In our notation, Theorem 3.7 states that for all graphs $G=(V,E)$ on $N$ vertices and any initial mutant location $u\in V$, we have $\mathsf{FT}_{r=1}^{\delta=1/2}(G,\{u\})\leq N^{5}/4$. In the case of $(d_{1},d_{2})$-bidegreed graphs, we are able to say more about their behavior relative to $\delta$. In particular, fixation times on bidegreed graphs are at most some polynomial in $N$, the degree ratio $d_{2}/d_{1}$, and $\delta$.
Theorem 3.8 (Fixation time on bidegreed graphs).
On a bidegreed graph with $d_{1}\leq d_{2}$, the fixation time of a neutral mutant starting on a vertex with degree $d_{1}$ is at most $N^{4}\cdot(d_{2}/d_{1})^{5}/4$. The fixation time of a neutral mutant starting on a vertex with degree $d_{2}$ is at most $N^{4}\cdot(d_{2}/d_{1})^{6}/4$.
For both Theorem 3.7 and Theorem 3.8, we also provide the absorption times starting from an arbitrary location and more precise bounds in the proofs in Appendices F.1 and F.2, respectively.
3.2 Constant selection
Not only do we obtain results for neutral evolution, but also for constant selection. Suppose mutants have relative fitness $r>1$ and wild-types have fitness $1$. We examine two regimes: weak selection ($r>1$ on any graph) and strong selection ($r>(D/d)^{2}$ where $D$ is the maximum degree and $d$ is the minimum degree of the graph). Despite the fact that fixation probabilities and times can be increasing, decreasing, or non-monotonic in $\delta$, we prove that nearly all graphs have short fixation times and provide an efficient algorithm to estimate fixation probabilities.
3.2.1 Weak selection
For undirected unweighted graph families $(G_{N})_{N\geq 1}$, it is known that for the pure Bd process with $r\neq 1$, $\mathsf{AT}_{r}^{\textrm{Bd}}(G_{N})$ is $o(N^{3+\varepsilon})$ for all $\varepsilon>0$ [49, 50]. For $r=1$, it is known $\mathsf{AT}_{r=1}^{\textrm{Bd}}(G_{N})$ is $O(N^{6})$ [1] and $\mathsf{AT}_{r=1}^{\textrm{dB}}(G_{N})$ is $O(N^{5})$ [48]. For any undirected unweighted graph $G=(V,E)$ on $N$ vertices when $r>1$ and $\delta=0,1$, we have $\mathsf{fp}_{r}^{\delta}(G)\geq\mathsf{fp}_{r=1}^{\delta}(G)=\Omega(N^{-2})$; thus $\mathsf{FT}_{r}^{\mathrm{Bd}}(G_{N})$ is $o(N^{5+\varepsilon})$ for all $\varepsilon>0$. Similar to fixation times in neutral evolution, we prove a polynomial bound for any mixed $\delta=1/2$ process.
Theorem 3.9 (Fixation time under weak selection).
If $\delta=1/2$ and $r>1$, then the fixation time of a mutant starting in an arbitrary location on any graph of size $N$ is at most $N^{5}\cdot\frac{r}{r-1}$.
In our notation, Theorem 3.9 states that for any graph $G=(V,E)$ on $N$ vertices and initial mutant location $u\in V$, we have $\mathsf{FT}_{r}^{\delta=1/2}(G,\{u\})\leq N^{5}\cdot r/(r-1)$. The proof is in Section 9 in the SI.
Figure 6: a, Fixation probabilities on an undirected cycle with $N=10$ vertices as a function of $r$ for $\delta=0,0.25,0.5,0.75,1$ starting with one mutant. The various curves represent the various values of $\delta$. The solid lines represent the average of $10^{4}$ simulations of the mixed $\delta$-updating process; the surrounding translucent regions are the $95\%$ confidence intervals. The dashed lines are the theoretical fixation probabilities from Theorem 3.10. i) zoomed in plot of the theoretical curves in the region $r\in[0.75,1.25]$, normalized by the fixation probability when $\delta=0.5$. At $r=1$, the fixation probability is $1/N$ regardless of $\delta$—this value is indicated by a dot.
For an undirected cycle $C_{N}$ on $N$ vertices, it is known for Bd that $\mathsf{fp}_{r}^{\mathrm{Bd}}(C_{N})=\mathsf{fp}_{r}^{\mathrm{Bd}}(K_{N})=\frac{1-1/r}{1-1/r^{N}}$ for all $r\in(0,1)\sqcup(1,\infty)$ and $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(C_{N})=1/N$ [34]. For dB, it is known [42, 6, 35] that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(C_{N})=1/N$, and for $r\in(0,1)\sqcup(1,\infty)$,
$\mathsf{fp}_{r}^{\mathrm{dB}}(C_{N})=\frac{2(r-1)}{3r-1+(r-3)r^{2-N}}.$ (2)
We prove a general formula for any $r$ and $\delta$.
Theorem 3.10 (Fixation probability on cycles).
Let $r>0$ and $\delta\in[0,1]$. Define $F_{k}:=rk+(N-k)$ to be the total fitness in the population with $k$ mutants. Define $\gamma_{k}$ as
[…]
Then
$\mathsf{fp}_{r}^{\delta}(C_{N})=\frac{1}{1+\sum_{j=1}^{N-1}\prod_{k=1}^{j}\gamma_{k}}.$ (4)
See Figure 6 for a plot of the fixation probability as a function of $r$ and $\delta$. See Section 10 in the SI for the proof.
Figure 7: Fixation probabilities on a star graph with $N=10$ vertices when $r\neq 1$. The various curves represent the various values of $\delta$. The solid lines represent the average of $10^{4}$ simulations of the mixed $\delta$-updating process; the surrounding translucent regions are the $95\%$ confidence intervals. The dashed lines are the theoretical fixation probabilities from Theorem 3.11. a, Fixation probabilities as a function of $r$ for $\delta=0,0.25,0.5,0.75,1$ starting with a mutant in the center. b, Same as a) except a mutant is initially on a leaf.
Let $S_{N}$ be a star graph with one center labelled $c$ and n:=N1n:=N-1 leaves labelled $\ell_{1},\ldots,\ell_{n}$. For Bd, it is known [4] that
$\mathsf{fp}_{r}^{\mathrm{Bd}}(S_{N},\{\ell_{i}\})=\frac{nr}{nr+1}\cdot p\quad\text{and}\quad\mathsf{fp}_{r}^{\mathrm{Bd}}(S_{N},\{c\})=\frac{r}{r+n}\cdot p$ (5)
with
$p=\frac{1}{1+\frac{n}{n+r}\sum_{j=1}^{n-1}\left(\frac{n+r}{r(nr+1)}\right)^{j}}.$ (6)
For dB, it is known [25, 35] that
$\mathsf{fp}_{r}^{\mathrm{dB}}(S_{N})=\frac{(N-1)r+1}{N(r+1)}\cdot\left(\frac{1}{N}+\frac{r}{N+2r-2}\right).$ (7)
We prove a general formula for any $r$ and $\delta$.
Theorem 3.11 (Fixation probability on stars).
For a star $S_{N}$ with center $c$ and n=N1n=N-1 leaves, for all $\delta\in[0,1]$ and $r>0$, the fixation probabilities $\mathsf{fp}_{r}^{\delta}(S_{N},\{c\})$ and $\mathsf{fp}_{r}^{\delta}(S_{N},\{\ell_{i}\})$ are given by explicit formulas provided in Section 11 in the SI. Furthermore, when $r=1$ (neutral evolution), these formulas recover Theorem 3.6.
See Figure 7 for a plot of the fixation probability as a function of $r$ and $\delta$ for a fixed star.
3.2.2 Strong selection
For pure Bd and dB, no known closed-form analytical formulas exist for all $r$. For Bd, there is a fully polynomial randomized approximation scheme (FPRAS) that with confidence $1-\alpha$ outputs an estimate of $\mathsf{fp}_{r}^{\mathrm{Bd}}(G)$ with multiplicative error $\varepsilon$ in time polynomial in $N$, $1/\varepsilon$, and $\log(1/\alpha)$ [1, 49, 51]. Furthermore, for both Bd and dB, there is a known polynomial-time (with respect to $N$) algorithm that computes $\partial\mathsf{fp}_{r}^{\delta}(G)/\partial r\big|_{r=1^{+}}$ for any undirected weighted graph $G$ [28]. For graphs with sufficiently large $r$ relative to the degree ratio, we prove the following bounds.
Theorem 3.12 (Fixation probability lower bound under strong selection).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$ and $r>(D/d)^{2}$, the fixation probability starting at an arbitrary location satisfies $\mathsf{fp}_{r}^{\delta}(G,\{u\})\geq 1/N^{2}$ for all $u\in V$.
Theorem 3.13 (Fixation time upper bound under strong selection).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$ and $r>(D/d)^{2}$, the fixation time starting at an arbitrary location is at most $N^{7}\cdot\frac{r}{r-1}$.
See Section 12 in the SI for the proofs of the preceding two theorems. Together, these two results show that when $r$ is large enough relative to the degree heterogeneity, the process behaves well. In particular, using the bounds from Theorems 3.12 and 3.13, we give a simple FPRAS for $\mathsf{fp}_{r}^{\delta}(G)$.
Theorem 3.14 (Fixation probability estimation algorithm).
For any graph $G=(V,E)$ on $N$ vertices with maximum degree $D$ and minimum degree $d$, for all $\delta\in[0,1]$, $r>(D/d)^{2}$, $\varepsilon>0$, and $\alpha\in(0,1]$, there exists a polynomial time (with respect to $N$, $1/\varepsilon$, and $\log(1/\alpha)$) randomized algorithm $\mathcal{A}$ that with probability at least $1-\alpha$ outputs an estimate of $\mathsf{fp}_{r}^{\delta}(G,S_{0})$ within multiplicative error $\varepsilon$.
See Section 15 in the SI for the proof. Note that when $\delta=1/2$, the bound on fixation time from Theorem 3.9 gives an FPRAS for all undirected unweighted graphs with $r>1$ (see Theorem 14.1 in the SI).
4 Conclusion
In this paper, we have studied mixed dB and Bd updating for neutral evolution and for constant selection. At each time step we use dB updating with probability $\delta$ and Bd updating with probability $1-\delta$. We have derived the following results that appear both in the main text and the appendices in full.
The fixation probability as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. Despite this, we find that the fixation probability is $1/N$ on any such graph when $\delta=1/2$. Additionally, it cannot be the case that the fixation probability is greater than $1/N$ for both pure dB and pure Bd (see Appendix D).
We give an explicit formula for the fixation probability on a bidegreed graph for any $\delta$. Further, we prove that with a deterministic initial mutant location, the fixation probability in neutral evolution on a bidegreed graph is the most sensitive at pure Bd or pure dB (see Appendix E.1).
Similar to the fixation probability, the fixation time as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. We prove that the expected (conditional) fixation time on such graphs is at most some polynomial in $N$ when $\delta=1/2$ (see Appendix F.1).
For constant selection with $r>1$, we prove that when $\delta=1/2$ the fixation time is at most $N^{5}\cdot r/(r-1)$ on any undirected unweighted graph (see Section 9 in the SI). We also prove that for graphs with $r>(D/d)^{2}$, the expected fixation time is at most $N^{7}\cdot r/(r-1)$ for any $\delta$ (see Section 12 in the SI). When $r>(D/d)^{2}$, we provide an efficient estimation algorithm for any $\delta$ for the fixation probability (see Section 15 in the SI).
For constant selection on undirected cycles and stars, we provide tractable analytical formulas for all $\delta$ (see Section 10 and Section 11 in the SI, respectively).
While fixation probabilities and times can have strange behaviors with respect to $\delta$, we have shown that unbiased mixing and nearly-regular graphs lead to well-behaved quantities.
There are many future directions of this work. We provide two promising directions. Firstly, one could classify the graphs that have monotone fixation probability in $\delta$, even in neutral evolution. Finally, one could estimate a realistic $\delta$ from data, by fitting the $\delta$-mixed updating model to empirical fixation probabilities in, for example, microbial populations, cancer cell lines, or laboratory evolution experiments.
Data Availability. Our code is available as a GitHub repository at https://github.com/davidb2/mixed-updating. For graph exploration, we utilized the computer program nauty [52]. We also explored dynamics using EvoLudo [53].
Acknowledgments. The authors thank Josef Tkadlec for many helpful suggestions. D.A.B. was supported by a Harvard Graduate School of Arts and Sciences Prize Fellowship. Y.H. and M.M. were supported by NSF grant CNS-2107078.
Disclaimer. A preliminary version of this work with differing content appears in an ArXiv preprint for the Innovations in Theoretical Computer Science 2026 conference [54].
Appendices
Appendix A Technical Tools
One main technical tool we will use is the martingale. We will use a potential function on the set $S\subseteq V$. The ultimate goal is to show that variance (i.e. second moment) of the mixed $1/2$-updating process as a function of the potential fluctuates a reasonable amount in expectation. We first recall the optional stopping theorems
Theorem A.1 (Optional Stopping Theorem for martingales, Corallary of Theorem 17.4 of [55]).
Let $k>0$ be some constant. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k\}\subseteq\mathcal{X}\subseteq[0,k]$. Let $\tau:=\min\{i:X_{i}\in\{0,k\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]=0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{E}[X_{\tau}]=\mathbb{E}[X_{0}]$.
Theorem A.2 (Optional Stopping Theorem for supermartingales, Corallary of Theorem 17.4 of [55]).
Let $k>0$ be some constant. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k\}\subseteq\mathcal{X}\subseteq[0,k]$. Let $\tau:=\min\{i:X_{i}\in\{0,k\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]\leq 0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{E}[X_{\tau}]\leq\mathbb{E}[X_{0}]$.
Corollary A.1 (Fixation probability of martingale).
Let $k_{2}\geq k_{1}>0$ be some constants. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k_{2}\}\subseteq\mathcal{X}\subseteq[0,k_{2}]$. Suppose $\mathbb{E}[X_{0}]=k_{1}$. Let $\tau:=\min\{i:X_{i}\in\{0,k_{2}\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]=0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{P}(X_{\tau}=k_{2})=k_{1}/k_{2}$.
Proof.
By Theorem A.2, $\mathbb{P}(X_{\tau}=k_{2})\cdot k_{2}=\mathbb{E}[X_{\tau}]=\mathbb{E}[X_{0}]=k_{1}$. ∎
Lemma A.1 (Absorption time is finite).
We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $1/2$-updating process where $Y_{i}\subseteq V$. Let $\phi\colon 2^{V}\to\mathbb{R}_{\geq 0}$ and $\tau:=\min\{i:\phi(Y_{i})\in\{0,\phi(V)\}\}$. Then $\mathbb{E}[\tau]<\infty$.
Proof.
Consider a contiguous block of $N$ successive states of the Markov chain $Y_{i},\dots,Y_{i+N-1}$ with $Y_{i}\not\in\{\emptyset,V\}$. Then it must be that $|Y_{i}|\geq 1$. Consider such a contiguous block in which some $u\in S_{i}$ spreads to the whole graph. Each step happens with probability at least $1/N^{2}$, so the entire sequence occurs with probability at least $p=N^{-2N}$. Thus, the expected number of such contiguous blocks until absorption occurs is at most $1/p<\infty$. ∎
Consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $\delta$-updating process where $Y_{i}\subseteq V$. Let $\phi\colon 2^{V}\to\mathbb{R}_{\geq 0}$. We define the following quantities for use in the proofs that follow:
$\psi_{u,v}^{\textrm{Bd}}(S)$ $:=\frac{1}{N}\left(\frac{1}{\deg(u)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{\deg(v)}(\phi(S)-\phi(S\setminus\{u\}))\right)$ (8)
$\psi_{u,v}^{\textrm{dB}}(S)$ $:=\frac{1}{N}\left(\frac{1}{\deg(v)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{\deg(u)}(\phi(S)-\phi(S\setminus\{u\}))\right)$ (9)
$\psi_{u,v}^{\delta}(S)$ $:=(1-\delta)\psi_{u,v}^{\textrm{Bd}}(S)+\delta\psi_{u,v}^{\textrm{dB}}(S).$ (10)
Then $\psi_{u,v}^{\delta}(S)=\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]$.
Appendix B Additivity of fixation probability
Lemma B.1 (Additivity in neutral evolution).
Let $G=(V,E)$ be a strongly connected weighted directed graph. For any $T\subseteq S\subseteq V$ we have
$\mathsf{fp}_{r=1}^{\delta}(G,S)=\mathsf{fp}_{r=1}^{\delta}(G,T)+\mathsf{fp}_{r=1}^{\delta}(G,S\setminus T).$ (11)
Proof.
Consider the variant of the mixed $\delta$-updating process where every vertex starts with a unique type. The process will evolve as usual until a single type takes over the whole graph. A mutant currently occupying subset $S$ fixates in the mixed $\delta$-updating process if and only if one of the vertices in the subset takes over in the altered process, which in turn equals the fixation probability with only that vertex by viewing all other types as wild types. ∎
Appendix C Proof of Theorem 3.2
Proof.
We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $1/2$-updating process where $Y_{i}\subseteq V$. Let $S\subseteq V$. We use the potential function $\phi(S)=|S|$. Let $\tau:=\min\{i:\phi(Y_{i})\in\{0,N\}\}$. By Lemma A.1, $\tau$ has finite expectation. Notice that $\psi_{u,v}^{\text{Bd}}(S)=-\psi_{u,v}^{\text{dB}}(S)$, which implies $\psi_{u,v}^{\delta=1/2}(S)=0$. Therefore $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]=\psi_{u,v}^{\delta=1/2}(S)=0$. Then we can apply Corollary A.1 to yield the desired result. ∎
Appendix D Proof that it is not possible for both pure processes to simultaneously have fixation probabilities greater than $1/N$
Lemma D.1.
Let $G=(V,E)$ and $u\in V$. Then it must be the case that
$\min\{\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\}),\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})\}\leq 1/N.$ (12)
Proof.
From [47, 44], we know that
$\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})\propto 1/\deg(u)\quad\text{and}\quad\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\})\propto\deg(u).$ (13)
Suppose that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\})>1/N$. Then it must be that $\deg(u)>D_{\textrm{AM}}$ where we take $D_{\textrm{AM}}:=\frac{1}{N}\cdot\sum_{v\in V}\deg(v)$} to be the arithmetic mean of the degrees of the vertices in $G$. Since the arithmetic mean is at least the harmonic mean of a sequence, we have that $\deg(u)\geq D_{\textrm{HM}}$ where $D_{\textrm{HM}}:=\frac{N}{\sum_{v\in V}1/\deg(v)}$} is the harmonic mean of the degrees of the vertices in $G$. Thus
$1/\deg(u)\leq 1/D_{\textrm{HM}}=\frac{1}{N}\sum_{v\in V}\frac{1}{\deg(v)}.$ (14)
This means that $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})\leq 1/N$.
On the other hand, suppose that $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})>1/N$. Then it must be that $\deg(u)<D_{\textrm{HM}}$. Since $D_{\textrm{HM}}\leq D_{\textrm{AM}}$, we have that $\deg(u)\leq D_{\textrm{AM}}$. Thus $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\})\leq 1/N$. ∎
Appendix E Proof of Theorem 3.5
Proof of Theorem 3.5.
We denote the function $B(d_{i})$ as $1$ if $i=1$ and $\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}$ if $i=2$. We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $\delta$-updating process where $Y_{i}\subseteq V$. We use the potential function $\phi\colon 2^{V}\to\mathbb{R}_{\geq 0}$ such that $\phi(S)=\sum_{v\in S}B(\deg(v))$. If $u\in S$ and $v\not\in S$, then
$\psi_{u,v}^{\delta}(S)=\frac{\delta}{N}\left(\frac{B(\deg(v))}{\deg(v)}-\frac{B(\deg(u))}{\deg(u)}\right)+\frac{1-\delta}{N}\left(\frac{B(\deg(v))}{\deg(u)}-\frac{B(\deg(u))}{\deg(v)}\right).$ (15)
  • If $\deg(u)=\deg(v)$, then clearly $\psi_{u,v}^{\delta}(S)=0$.
  • Otherwise, if $\deg(u)<\deg(v)$ then
    $\psi_{u,v}^{\delta}(S)$ $=\frac{1}{N}\cdot\left[\delta\left(\frac{B(d_{2})}{d_{2}}-\frac{1}{d_{1}}\right)+(1-\delta)\left(\frac{B(d_{2})}{d_{1}}-\frac{1}{d_{2}}\right)\right]$ (16)
    $=\frac{1}{Nd_{1}d_{2}}\cdot\left[\delta(B(d_{2})d_{1}-d_{2})+(1-\delta)(B(d_{2})d_{2}-d_{1})\right]$ (17)
    $=\frac{1}{Nd_{1}d_{2}}\cdot\left[B(d_{2})((1-\delta)d_{2}+\delta d_{1})-((1-\delta)d_{1}+\delta d_{2})\right]$ (18)
    $=0.$ (19)
  • Due to symmetry, the case of $\deg(u)>\deg(v)$ follows from the same argument as the preceeding case.
Thus in any case, $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]=0$. Applying Corollary A.1 with $k_{1}=B(\deg(u))$ and $k_{2}=\sum_{v\in V}B(\deg(v))$ completes the proof. ∎
E.1 Sensitivity of fixation probabilities on bidegreed graphs
Theorem E.1.
Let $G=(V,E)$ be a $(d_{1},d_{2})$-bidegreed graph on $N\geq 1$ vertices with $d_{1}\leq d_{2}$. Let $n_{1}$ be the number of vertices of degree $d_{1}$ and $n_{2}$ the number of vertices of degree $d_{2}$ such that $N=n_{1}+n_{2}$ is the total number of vertices. For $\delta\in[0,1]$, let
$\rho_{1}(\delta):=\mathsf{fp}_{r=1}^{\delta}(G,\{u\})\quad\text{for any }u\in V\text{ with }\deg(u)=d_{1},$ (20)
$\rho_{2}(\delta):=\mathsf{fp}_{r=1}^{\delta}(G,\{u\})\quad\text{for any }u\in V\text{ with }\deg(u)=d_{2}.$ (21)
Then for all $\delta\in[0,1]$, the values $\rho^{\prime}_{1}(\delta)$ and $\rho^{\prime}_{2}(\delta)$ exist and are given by $\rho_{1}^{\prime}(\delta)=-n_{2}\cdot Q(\delta)$ and $\rho_{2}^{\prime}(\delta)=n_{1}\cdot Q(\delta)$ where
$Q(\delta):=\frac{(d_{2}-d_{1})(d_{1}+d_{2})}{D(\delta)^{2}}$ (22)
and
$D(\delta):=d_{1}n_{2}+d_{2}n_{1}+\delta(d_{1}-d_{2})(n_{1}-n_{2}).$ (23)
If $d_{1}<d_{2}$, then for all $\delta\in[0,1]$ one has $\rho_{1}^{\prime}(\delta)<0$ and $\rho_{2}^{\prime}(\delta)>0$. Otherwise, $\rho_{1}^{\prime}(\delta)=\rho_{2}^{\prime}(\delta)=0$ for all $\delta$.
Proof.
We denote the function $B(d_{i})$ as $1$ if $i=1$ and $\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}$ if $i=2$. By Theorem 3.5, for any vertex $u\in V$ with degree $d_{1}$ or $d_{2}$ we have
$\mathsf{fp}_{r=1}^{\delta}(G,\{u\})=\frac{B(\deg(u))}{\sum_{v\in V}B(\deg(v))}.$ (24)
Let
$S(\delta):=\sum_{v\in V}B(\deg(v))=n_{1}+n_{2}B_{2}(\delta)\quad\text{where}\quad B_{2}(\delta):=\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}.$ (25)
Then $\rho_{1}(\delta)=\frac{1}{S(\delta)}$ and $\rho_{2}(\delta)=\frac{B_{2}(\delta)}{S(\delta)}$.
We first compute $B_{2}^{\prime}(\delta)$. Write $B_{2}(\delta)=A(\delta)/C(\delta)$ where
$A(\delta)=(1-\delta)d_{1}+\delta d_{2}=d_{1}+\delta(d_{2}-d_{1}),$ (26)
$C(\delta)=(1-\delta)d_{2}+\delta d_{1}=d_{2}+\delta(d_{1}-d_{2}).$ (27)
Then $A^{\prime}(\delta)=d_{2}-d_{1}$ and $C^{\prime}(\delta)=d_{1}-d_{2}$. By the quotient rule,
$B_{2}^{\prime}(\delta)=\frac{A^{\prime}C-AC^{\prime}}{C^{2}}=\frac{(d_{2}-d_{1})C-(d_{1}+d_{2}+\delta(d_{2}-d_{1}))(d_{1}-d_{2})}{C^{2}}.$ (28)
A short algebraic simplification yields
$B_{2}^{\prime}(\delta)=\frac{d_{2}^{2}-d_{1}^{2}}{C(\delta)^{2}}=\frac{(d_{2}-d_{1})(d_{2}+d_{1})}{\bigl((1-\delta)d_{2}+\delta d_{1}\bigr)^{2}}.$ (29)
In particular $B_{2}^{\prime}(\delta)$ has the same sign as $(d_{2}-d_{1})$.
Next we differentiate $S(\delta)$ to get $S^{\prime}(\delta)=n_{2}B_{2}^{\prime}(\delta)$. Hence
$\rho_{1}^{\prime}(\delta)=-\frac{S^{\prime}(\delta)}{S(\delta)^{2}}=-\frac{n_{2}B_{2}^{\prime}(\delta)}{S(\delta)^{2}}.$ (30)
Substituting the explicit form of $B_{2}^{\prime}(\delta)$ and clearing denominators, one finds
$\rho_{1}^{\prime}(\delta)=-\frac{n_{2}(d_{2}-d_{1})(d_{1}+d_{2})}{D(\delta)^{2}},$ (31)
One can check directly that $D(\delta)$ equals $S(\delta)C(\delta)$ up to a positive constant factor, so $D(\delta)\neq 0$ for all $\delta\in[0,1]$ and the sign of $\rho_{1}^{\prime}(\delta)$ agrees with the sign of $(d_{1}-d_{2})$ which is non-positive.
For $\rho_{2}$, differentiating $\rho_{2}(\delta)=B_{2}(\delta)/S(\delta)$ and using the quotient rule gives
$\rho_{2}^{\prime}(\delta)=\frac{B_{2}^{\prime}S-B_{2}S^{\prime}}{S^{2}}=\frac{B_{2}^{\prime}(\delta)\bigl(S(\delta)-n_{2}B_{2}(\delta)\bigr)}{S(\delta)^{2}}=\frac{B_{2}^{\prime}(\delta)n_{1}}{S(\delta)^{2}}.$ (32)
Substituting $B_{2}^{\prime}(\delta)$ again and simplifying yields
$\rho_{2}^{\prime}(\delta)=\frac{n_{1}(d_{2}-d_{1})(d_{1}+d_{2})}{D(\delta)^{2}}.$ (33)
Thus when $d_{1}<d_{2}$ we have $\rho_{1}^{\prime}(\delta)<0$ and $\rho_{2}^{\prime}(\delta)>0$ for all $\delta$. Therefore $\rho_{1}$ is strictly decreasing and $\rho_{2}$ strictly increasing on $[0,1]$ when $d_{1}<d_{2}$, and the endpoint maximum and minimum follow immediately.
When $d_{1}=d_{2}$, we get that $\rho_{1}^{\prime}(\delta)=\rho_{2}^{\prime}(\delta)$ for all $\delta\in[0,1]$. Therefore $\rho_{1}$ and $\rho_{2}$ are constant for all $\delta\in[0,1]$.
For the magnitudes $|\rho_{1}^{\prime}(\delta)|$ and $|\rho_{2}^{\prime}(\delta)|$ note that by Theorem E.1,
$|\rho_{1}^{\prime}(\delta)|=\frac{n_{2}(d_{2}-d_{1})(d_{1}+d_{2})}{D(\delta)^{2}},\qquad|\rho_{2}^{\prime}(\delta)|=\frac{n_{1}(d_{2}-d_{1})(d_{1}+d_{2})}{D(\delta)^{2}},$ (34)
so the $\delta$-dependence is entirely in the factor $1/D(\delta)^{2}$. Notice that $D$ is strictly monotone in $\delta$ unless $n_{1}=n_{2}$ or $d_{1}=d_{2}$, in which case $D$ is constant. If $n_{1}>n_{2}$ then $n_{1}-n_{2}>0$ and $d_{1}-d_{2}<0$, so $D$ is strictly decreasing on $[0,1]$, hence $1/D(\delta)^{2}$ and therefore $|\rho_{i}^{\prime}(\delta)|$ is strictly increasing in $\delta$, so its maximum occurs at $\delta=1$ and minimum at $\delta=0$. If $n_{1}<n_{2}$ then $n_{1}-n_{2}<0$ and $D$ is strictly increasing, so $|\rho_{i}^{\prime}(\delta)|$ is strictly decreasing, with maximum at $\delta=0$ and minimum at $\delta=1$. If $n_{1}=n_{2}$ then $D(\delta)$ is independent of $\delta$ and hence $|\rho_{i}^{\prime}(\delta)|$ is constant in $\delta$.
Appendix F Fixation time
F.1 Proof of Theorem 3.7
Proof.
We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $1/2$-updating process where $Y_{i}\subseteq V$. Let $\phi(S)=|S|$ and $\tau:=\min\{i:\phi(Y_{i})\in\{0,N\}\}$. We want to use Theorem A.2. By Lemma A.1, $\tau$ has finite expectation.
We now bound $\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]$ for all $S\notin\{\emptyset,V\}$. For any $S\notin\{\emptyset,V\}$ and edge $(u,v)\in E$ such that either (i) $u\in S$ and $v\not\in S$ or (ii) $u\not\in S$ and $v\in S$, with probability $1/N^{2}$ the step happens along $(u,v)$ and $\phi$ changes by $\pm 1$. So $\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]\geq 1/N^{2}$.
Define g(x)=x(Nx)g(x)=x(N-x) and $Z_{i}=g(\phi(Y_{i}))+\frac{i}{N^{2}}$. For any $x$ and $\Delta\in\{-1,0,1\}$,
$g(x+\Delta)-g(x)=(N-2x)\Delta-\Delta^{2}.$ (35)
Notice that $\psi_{u,v}^{\text{Bd}}(S)=-\psi_{u,v}^{\text{dB}}(S)$, which implies $\psi_{u,v}^{\delta=1/2}(S)=0$. Therefore $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]=\psi_{u,v}^{\delta=1/2}(S)=0$, and we get
$\mathbb{E}[Z_{i+1}-Z_{i}\mid Y_{i}=S]$ $=\mathbb{E}[g(\phi(Y_{i+1}))-g(\phi(Y_{i}))\mid Y_{i}=S]+\frac{1}{N^{2}}$ (36)
$=-\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]+\frac{1}{N^{2}}\leq 0,$ (37)
so $(Z_{i})$ is a supermartingale with bounded one-step differences. Thus, applying Theorem A.2 at the absorption time $\tau$, we get $\mathbb{E}[Z_{\tau}]\leq\mathbb{E}[Z_{0}]=g(\phi(Y_{0}))$. Since $\phi(Y_{\tau})\in\{0,N\}$, we have $g\left(\phi(Y_{\tau})\right)=0$ and $g(\phi(Y_{0}))=\phi(Y_{0})(N-\phi(Y_{0}))\leq N^{2}/4$. Thus when $Y_{0}=S_{0}$,
$\mathsf{AT}_{r=1}^{\delta=1/2}(G,S_{0})=\mathbb{E}[\tau]\leq N^{2}\cdot g\left(\phi(S_{0})\right)\leq\frac{N^{4}}{4}.\qed$ (38)
The fixation time bound follows from Theorem 3.2.
F.2 Proof of Theorem 3.8
Proof.
The proof is similar to that of Theorem 3.7 but more general. The only difference is that we use a different potential function and can bound the expected squared potential change more carefully using the bidegreed property.
We denote the function $B(d_{i})$ as $1$ if $i=1$ and $\frac{(1-\delta)d_{1}+\delta d_{2}}{(1-\delta)d_{2}+\delta d_{1}}$ if $i=2$. Let $S(\delta):=\sum_{u\in V}B(\deg(u))=n_{1}+n_{2}B(d_{2})$. We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $\delta$-updating process where $Y_{i}\subseteq V$. Let $\phi(S)=\sum_{u\in S}B(\deg(u))$ and $\tau:=\min\{i:\phi(Y_{i})\in\{0,\phi(V)\}\}$. By Lemma A.1, $\tau$ has finite expectation.
We now bound $\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]$ for all $S\notin\{\emptyset,V\}$. For any $S\notin\{\emptyset,V\}$ and edge $(u,v)\in E$ such that either (i) $u\in S$ and $v\not\in S$ or (ii) $u\not\in S$ and $v\in S$, with probability $1/N^{2}$ the step happens along $(u,v)$ and $\phi$ changes by at most $d_{1}/d_{2}$ in either direction. So $\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]\geq\left(\frac{d_{1}/d_{2}}{N}\right)^{2}$.
Define $g(x)=x(\phi(V)-x)$ and $Z_{i}=g(\phi(Y_{i}))+i\cdot\left(\frac{d_{1}/d_{2}}{N}\right)^{2}$. For $\Delta_{i}:=\phi(Y_{i+1})-\phi(Y_{i})$,
$g(\Delta_{i})=(\phi(V)-2\phi(Y_{i}))\cdot\Delta_{i}-\Delta_{i}^{2}.$ (39)
Therefore using the fact from the proof of Theorem 3.5 that $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]=0$, we get
$\mathbb{E}[Z_{i+1}-Z_{i}\mid Y_{i}=S]$ $=\mathbb{E}[g(\phi(Y_{i+1}))-g(\phi(Y_{i}))\mid S_{i}]+\left(\frac{d_{1}/d_{2}}{N}\right)^{2}\$(40)
$=-\mathbb{E}[(\phi(Y_{i+1})-\phi(Y_{i}))^{2}\mid Y_{i}=S]+\left(\frac{d_{1}/d_{2}}{N}\right)^{2}\leq 0,$ (41)
so $(Z_{i})$ is a supermartingale with bounded one-step differences. Thus, applying Theorem A.2 at the absorption time $\tau$, we get $\mathbb{E}[Z_{\tau}]\leq\mathbb{E}[Z_{0}]=g(\phi(Y_{0}))$. Since $\phi(Y_{\tau})\in\{0,\phi(V)\}$, we have $g\left(\phi(Y_{\tau})\right)=0$ and $g(\phi(Y_{0}))=\phi(Y_{0})(\phi(V)-\phi(Y_{0}))\leq\phi(V)^{2}/4$. Let $S(\delta)=\phi(V)$. Thus when $Y_{0}=S_{0}$,
$\mathsf{AT}_{r=1}^{\delta}(G,S_{0})=\mathbb{E}[\tau]\leq N^{2}\cdot(d_{2}/d_{1})^{2}\cdot g\left(\phi(S_{0})\right)\leq\frac{S(\delta)^{2}\cdot(d_{2}/d_{1})^{2}\cdot N^{2}}{4}.\qed$ (42)
The fixation time bound follows from applying the fixation probability formula from Theorem 3.5 and the fact that $S(\delta)\leq N\cdot(d_{2}/d_{1})$.
Supplementary Information
7 Proof of Theorem 3.6
Proof.
Let $d_{i}=n-i$ denote the number of wild-type leaves. At $r=1$,
$F_{i}^{\star}=F_{i}^{\varnothing}=n+1,\qquad G_{i}=n,$ (43)
so the coefficients from above simplify to
$A_{i}$ $=(1-\delta)\frac{d_{i}}{n(n+1)}+\delta\frac{d_{i}}{n+1}=\frac{n-i}{n+1}\Big(\frac{(1-\delta)}{n}+\delta\Big),$ (44)
$B_{i}$ $=(1-\delta)\frac{d_{i}}{n+1}+\delta\frac{d_{i}}{n(n+1)}=\frac{n-i}{n+1}\Big((1-\delta)+\frac{\delta}{n}\Big),$ (45)
$C_{i}$ $=\frac{i}{n},$ (46)
$a_{i}$ $=(1-\delta)\frac{i}{n+1}+\delta\frac{i}{n(n+1)}=\frac{i}{n+1}\Big((1-\delta)+\frac{\delta}{n}\Big),$ (47)
$c_{i}$ $=(1-\delta)\frac{i}{n(n+1)}+\delta\frac{i}{n+1}=\frac{i}{n+1}\Big(\frac{(1-\delta)}{n}+\delta\Big),$ (48)
$b_{i}$ $=1-a_{i}-c_{i}=1-\frac{i}{n}.$ (49)
Normalizing gives
$\alpha_{i}=\frac{A_{i}}{1-C_{i}},\quad\beta_{i}=\frac{B_{i}}{1-C_{i}},\quad p_{i}=\frac{a_{i}}{1-b_{i}},\quad q_{i}=\frac{c_{i}}{1-b_{i}}.$ (50)
Since $1-C_{i}=(n-i)/n$ and $1-b_{i}=i/n$, the factors (ni)(n-i) and $i$ cancel, and all four quantities become
$\alpha=\frac{n\delta+(1-\delta)}{n+1},\qquad\beta=\frac{n(1-\delta)+\delta}{n+1},\qquad p=\beta,\qquad q=\alpha,\qquad\alpha+\beta=1.$ (51)
With these constants, the two recurrences read
$P_{i}^{\star}=\alpha P_{i+1}^{\star}+\beta P_{i}^{\varnothing},\qquad P_{i}^{\varnothing}=\beta P_{i}^{\star}+\alpha P_{i-1}^{\varnothing}.$ (52)
This can be written in $2\times 2$ form as
$\begin{pmatrix}P_{i+1}^{\star}\\ P_{i}^{\varnothing}\end{pmatrix}=M\begin{pmatrix}P_{i}^{\star}\\ P_{i-1}^{\varnothing}\end{pmatrix},\qquad M=\begin{pmatrix}\frac{1-\beta p}{\alpha}&-\frac{\beta q}{\alpha}\\[2.0pt] p&q\end{pmatrix}=\begin{pmatrix}1+\beta&-\beta\\ \beta&\alpha\end{pmatrix}$. (53)
We will raise this matrix to a power. Note that
$\mathrm{tr}(M)=(1+\beta)+\alpha=2,\qquad\det(M)=(1+\beta)\alpha-(-\beta)\beta=(1+\beta)(1-\beta)+\beta^{2}=1,$ (54)
so $M$ has eigenvalue $1$ with algebraic multiplicity $2$. Indeed,
$M-I=\begin{pmatrix}\beta&-\beta\\ \beta&-\beta\end{pmatrix}=\beta\begin{pmatrix}1&-1\\ 1&-1\end{pmatrix}{=:Q},\qquad Q^{2}=0,$ (55)
hence $(M-I)^{2}=0$. Therefore by the binomial theorem, for every $t\in\mathbb{Z}_{\geq 0}$,
$M^{t}=(Q+I)^{t}=I+tQ=\begin{pmatrix}1+t\beta&-t\beta\\ t\beta&1-t\beta\end{pmatrix}$. (56)
Now apply the boundary conditions $P_{0}^{\varnothing}=0$ and $P_{n}^{\star}=1$. Iterating the relation $(P_{i+1}^{\star},P_{i}^{\varnothing})^{\top}=M(P_{i}^{\star},P_{i-1}^{\varnothing})^{\top}$ gives
$\begin{pmatrix}P_{n}^{\star}\\ P_{n-1}^{\varnothing}\end{pmatrix}=M^{n-1}\begin{pmatrix}P_{1}^{\star}\\ P_{0}^{\varnothing}\end{pmatrix}=M^{n-1}\begin{pmatrix}P_{1}^{\star}\\ 0\end{pmatrix}$. (57)
Taking the first coordinate and using (56),
$1=P_{n}^{\star}=(M^{n-1})_{11}P_{1}^{\star}=\big(1+(n-1)\beta\big)P_{1}^{\star},$ (58)
so
$P_{1}^{\star}=\frac{1}{1+(n-1)\beta}.$ (59)
The identities $P_{0}^{\star}=\alpha P_{1}^{\star}$ and $P_{1}^{\varnothing}=pP_{1}^{\star}=\beta P_{1}^{\star}$ then give
$P_{0}^{\star}=\frac{\alpha}{\alpha+n\beta},\qquad P_{1}^{\varnothing}=\frac{\beta}{\alpha+n\beta}.$ (60)
It is convenient to parametrize by $z:=\alpha/\beta$ (well-defined since $\beta>0$ for $n\geq 1$). Then
$P_{0}^{\star}=\frac{z}{n+z},\qquad P_{1}^{\varnothing}=\frac{1}{n+z}.$ (61)
A short calculation with the explicit $\alpha,\beta$ yields
$z=\frac{\alpha}{\beta}=\frac{n\delta+(1-\delta)}{(1-\delta)(n-1)+1},$ (62)
and the general formulas
$P_{i}^{\varnothing}=\frac{i}{n+z},\qquad P_{i}^{\star}=\frac{i+z}{n+z}$ (63)
follow by a single-step induction using the two recurrences (they also match the boundary values $P_{0}^{\varnothing}=0,P_{n}^{\star}=1$). Replacing n=N1n=N-1 gives the formulas in $N$. ∎
8 Technical Tools for Constant Selection
The following theorem is proven in [1].
Theorem 8.1 (Absorption time, negative drift).
Let $(Y_{i})_{i\geq 0}$ be a Markov chain with state space $\Omega$, where $Y_{0}$ is chosen from some set $I\subseteq\Omega$. If there are constants $k_{1},k_{2}>0$ and a non-negative function $\phi:\Omega\to\mathbb{R}$ such that
  • $\phi(S)=0$ for some $S\in\Omega$,
  • $\phi(S)\leq k_{1}$ for all $S\in I$, and
  • $\mathbb{E}[\phi(Y_{i})-\phi(Y_{i+1})\mid Y_{i}=S]\geq k_{2}$ for all $i\geq 0$ and all $S$ with $\phi(S)>0$,
then $\mathbb{E}[\tau]\leq k_{1}/k_{2}$, where $\tau=\min\{i:\phi(Y_{i})=0\}$.
Corollary 8.1 (Absorption time, positive drift).
Let $(Y_{i})_{i\geq 0}$ be a Markov chain with state space $\Omega$, where $Y_{0}$ is chosen from some set $I\subseteq\Omega$. If there are constants $k_{1},k_{2}>0$ and a non-negative function $\phi:\Omega\to\mathbb{R}$ such that
  • $\phi(S)=0$ for some $S\in\Omega$, and $\phi(S)=k_{1}$ for some $S\in\Omega$,
  • $\phi(S)\leq k_{1}$ for all $S\in I$, and
  • $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]\geq k_{2}$ for all $i\geq 0$ and all $S$ with $\phi(S)\notin\{0,k_{1}\}$,
then $\mathbb{E}[\tau]\leq k_{1}/k_{2}$, where $\tau=\min\{i:\phi(Y_{i})=0\vee\phi(Y_{i})=k_{1}\}$.
Proof.
Consider the process $Y_{i}^{\prime}$ which behaves identically to $Y_{i}$ except that if $\phi(Y^{\prime}_{i})=0$, then $\phi(Y^{\prime}_{i+1})=S$ for some $S$ with $\phi(S)=k_{1}$. Let $\phi^{\prime}(S)=k_{1}-\phi(S)$. Then $\min\{i:\phi^{\prime}(Y_{i}^{\prime})=0\}=\min\{i:\phi(Y_{i}^{\prime})=k_{1}\}\geq\tau$. Now $\phi^{\prime}(Y_{i}^{\prime})$ satisfies the conditions in Theorem 8.1 with the same parameters. Applying Theorem 8.1 gives the claim. ∎
Theorem 8.2 (Fixation probability, lower bound).
Let $(Y_{i})_{i\geq 0}$ be a Markov chain with state space $\Omega$, where $Y_{0}$ is chosen from some set $I\subseteq\Omega$. Suppose there are constants $k_{2}\geq k_{1}>0$ and a non-negative function $\phi\colon\Omega\to\mathbb{R}$ such that
  • $\phi(S)=0$ for some $S\in\Omega$, and $\phi(S)=k_{2}$ for some $S\in\Omega$,
  • $\phi(Y_{0})\geq k_{1}$,
  • $\phi(S)\leq k_{2}$ for all $S\in I$,
  • $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]\geq 0$ for all $i\geq 0$ and all $S$ with $\phi(S)\notin\{0,k_{2}\}$.
Then $\mathbb{P}(\phi(Y_{\tau})=k_{2})\geq k_{1}/k_{2}$, where $\tau=\min\{i\geq 0:\phi(Y_{i})\in\{0,k_{2}\}\}$.
Proof.
By the claimed properties, $(\phi(Y_{i}))_{i\geq 0}$ is a submartingale. By the optional stopping theorem, $\mathbb{P}(\phi(Y_{\tau})=k_{2})\cdot k_{2}=\mathbb{E}[\phi(Y_{\tau})]\geq\mathbb{E}[\phi(Y_{0})]\geq k_{1}$. ∎
A particular notion will be useful in our proofs, describing the expected change due to a specific edge $(u,v)$:
$\psi_{u,v}^{\textrm{Bd}}(S)$ $=\frac{1}{w(S)}\left(\frac{r}{\deg(u)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{\deg(v)}(\phi(S)-\phi(S\setminus\{u\}))\right),$ (64)
$\psi_{u,v}^{\textrm{dB}}(S)$ $=\frac{1}{N}\left(\frac{r}{w_{v}(S)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{w_{u}(S)}(\phi(S)-\phi(S\setminus\{u\}))\right),$ (65)
$\psi_{u,v}^{\delta}(S)$ $=(1-\delta)\psi_{u,v}^{\textrm{Bd}}(S)+\delta\psi_{u,v}^{\textrm{dB}}(S),$ (66)
where w(S)=N+(r1)|S|w(S)=N+(r-1)|S| is the total fitness and $w_{u}(S)=\sum_{v:(u,v)\in E}f_{S}(v)$ is the fitness sum of neighbors of $u$.
9 Proof of Theorem 3.9
Theorem 9.1.
For any undirected unweighted graph $G=(V,E)$ on $N$ vertices and for all $r>1$, with $\phi(S)=|S|$ we have
$\psi^{\delta=1/2}_{u,v}(S)\geq\frac{r-1}{rN^{3}}$ (67)
for all $u,v\in V$, and
$\mathsf{AT}^{\delta=1/2}_{r}(G)\leq\frac{r}{r-1}\cdot N^{4}.$ (68)
Proof.
The potential function is $\phi(S)=|S|$. Thus
$\psi_{u,v}^{\delta=1/2}(S)$ $=\frac{1}{2}\left[\frac{1}{w(S)}\left(\frac{r}{\deg(u)}-\frac{1}{\deg(v)}\right)+\frac{1}{N}\left(\frac{r}{w_{v}(S)}-\frac{1}{w_{u}(S)}\right)\right].$ (69)
For all $S\neq V$ and $S\neq\emptyset$, we have $N-1+r\leq w(S)\leq(N-1)r+1$, therefore
$\frac{1}{w(S)}\frac{r}{\deg(u)}-\frac{1}{N}\frac{1}{w_{u}(S)}$ $\geq\frac{r}{(r(N-1)+1)\deg(u)}-\frac{1}{N\deg(u)}=\frac{1}{\deg(u)}\frac{1}{N}\frac{r-1}{Nr-r+1}\geq\frac{r-1}{rN^{3}}.$ (70)
Similarly,
$\frac{1}{N}\frac{r}{w_{v}(S)}-\frac{1}{w(S)}\frac{1}{\deg(v)}$ $\geq\frac{r}{rN\deg(v)}-\frac{1}{(N-1+r)\deg(v)}=\frac{1}{\deg(v)}\frac{1}{N}\frac{r-1}{N+r-1}\geq\frac{r-1}{rN^{3}}.$ (71)
Thus $\psi^{\delta=1/2}_{u,v}(S)\geq\frac{r-1}{rN^{3}}$, and the absorption time bound follows from Corollary 8.1. ∎
Lemma 9.1 (Coupling with fitnesses $1$ and $r$ at fixed $\delta$).
Fix an undirected unweighted graph $G=(V,E)$ on $N$ vertices and let $\delta\in[0,1]$. There exists a coupling of two mixed $\delta$-updating chains, $S_{t}^{(1)}$ with fitness $r_{1}=1$ and $S_{t}^{(2)}$ with fitness $r_{2}\geq r_{1}$, starting from the same $S_{0}\subseteq V$, such that $S_{t}^{(2)}\supseteq S_{t}^{(1)}$ for all $t\geq 0$.
Proof.
The claim holds at $t=0$. Assume inductively that $S_{\ell}^{(2)}\supseteq S_{\ell}^{(1)}$ for all $\ell\leq t$. At step $t+1$, draw $D\sim\mathrm{Bernoulli}(\delta)$.
Bd case ($D=0$). Draw $U\sim\mathrm{Unif}(0,1)$. Let $p_{i}(S_{t}^{(i)})=\frac{r_{i}|S_{t}^{(i)}|}{r_{i}|S_{t}^{(i)}|+N-|S_{t}^{(i)}|}$ be the probability that the parent is mutant in chain $i$. Since $r_{2}\geq r_{1}=1$, we have $p_{2}(S_{t}^{(2)})\geq p_{1}(S_{t}^{(1)})$ by the inductive hypothesis. Let the birth parent be mutant in chain $i$ when $U\leq p_{i}(S_{t}^{(i)})$. If the parent types agree, pick the same specific parent and target neighbor; then $S_{t+1}^{(2)}\supseteq S_{t+1}^{(1)}$. If chain $1$ selects wild-type and chain $2$ selects mutant, pick death targets independently; $S_{t+1}^{(2)}\supseteq S_{t+1}^{(1)}$ since we can only lose a mutant in chain $1$ or gain one in chain $2$.
dB case ($D=1$). Draw a common death vertex $v$ uniformly from $V$. Let $q_{i}=\frac{r_{i}\deg_{S_{t}^{(i)}}(v)}{r_{i}\deg_{S_{t}^{(i)}}(v)+\deg_{V\setminus S_{t}^{(i)}}(v)}$ be the probability the replacement is mutant in chain $i$. Since $r_{2}\geq r_{1}=1$, we have $q_{2}\geq q_{1}$ by the inductive hypothesis. Draw $U^{\prime}\sim\mathrm{Unif}(0,1)$ and declare a mutant parent in chain $i$ when $U^{\prime}\leq q_{i}$. The same argument as the Bd case gives $S_{t+1}^{(2)}\supseteq S_{t+1}^{(1)}$. ∎
Corollary 9.1.
For an undirected unweighted graph $G=(V,E)$ with $N$ vertices and $S_{0}\subseteq V$ with $r>1$, we have $\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\geq|S_{0}|/N$.
Proof.
By Lemma 9.1, for any $t$, $S_{t}^{(2)}\supseteq S_{t}^{(1)}$. Since $\mathsf{fp}_{r=1}^{\delta=1/2}(G,S_{0})=|S_{0}|/N$ (by Theorem 3.2), and fixation of chain $1$ implies fixation of chain $2$, the claim follows. ∎
Proof of Theorem 3.9.
We know $\mathsf{FT}_{r}^{\delta}(G,S_{0})\leq\mathsf{AT}_{r}^{\delta}(G,S_{0})/\mathsf{fp}_{r}^{\delta}(G,S_{0})$. By Theorem 9.1, $\mathsf{AT}_{r}^{\delta=1/2}(G,S_{0})\leq\frac{r}{r-1}\cdot N^{4}$. By Corollary 9.1, $\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\geq|S_{0}|/N\geq 1/N$. Thus $\mathsf{FT}_{r}^{\delta=1/2}(G,\{u\})\leq\frac{r}{r-1}\cdot N^{4}\cdot N=\frac{r}{r-1}\cdot N^{5}$. ∎
10 Proof of Theorem 3.10
We consider the fixation probability on a cycle $C_{N}$ with $N$ vertices starting with one mutant. Since the current number of mutants completely determines the state (mutants are all adjacent), we only need to compute the ratio $\gamma_{k}$ of the probability of decreasing over the probability of increasing when there are $k$ mutants. From this, the standard formula (see [32]) gives
$\mathsf{fp}_{r}^{\delta}(C_{N})=\frac{1}{1+\sum_{j=1}^{N-1}\prod_{k=1}^{j}\gamma_{k}}.$ (72)
Let $F_{k}:=rk+(N-k)$ be the total fitness of the population with $k$ mutants. The probability of increasing the mutant count with k<N1k<N-1 mutants is
$p_{k}^{\uparrow}=(1-\delta)r/F_{k}+2r\delta/((1+r)N).$ (73)
When k=N1k=N-1,
$p_{N-1}^{\uparrow}=(1-\delta)r/F_{N-1}+\delta/N.$ (74)
The probability of decreasing with $k>1$ mutants is
$p_{k}^{\downarrow}=(1-\delta)/F_{k}+2\delta/((1+r)N).$ (75)
When $k=1$,
$p_{1}^{\downarrow}=(1-\delta)/F_{1}+\delta/N.$ (76)
Thus
$\gamma_{k}=p_{k}^{\downarrow}/p_{k}^{\uparrow}=\begin{cases}\displaystyle\frac{(1-\delta)/F_{1}+\delta/N}{(1-\delta)r/F_{1}+2r\delta/((1+r)N)}&\text{if }k=1,\\[10.0pt] \displaystyle\frac{(1-\delta)/F_{N-1}+2\delta/((1+r)N)}{(1-\delta)r/F_{N-1}+\delta/N}&\text{if }k=N-1,\\[10.0pt] \displaystyle\frac{(1-\delta)/F_{k}+2\delta/((1+r)N)}{(1-\delta)r/F_{k}+2r\delta/((1+r)N)}&\text{otherwise.}\end{cases}$ (77)
11 Proof of Theorem 3.11
We derive an explicit formula for a star graph for general $r>0$ and $\delta\in[0,1]$. We consider a star graph $S_{N}$ with a center vertex labelled $c$ and n=N1n=N-1 leaves labelled $\ell_{1},\ldots,\ell_{n}$. A state in our Markov chain is $(i,\sigma)$ where $i\in\{0,\dots,n\}$ counts mutant leaves and $\sigma\in\{\star,\varnothing\}$ records whether the center is mutant ($\star$) or wild-type ($\varnothing$). We write
$P_{i}^{\star}:=\mathbb{P}\{\text{fixation}\mid(i,\star)\},\qquad P_{i}^{\varnothing}:=\mathbb{P}\{\text{fixation}\mid(i,\varnothing)\}.$ (78)
Boundary conditions: $P_{0}^{\varnothing}=0$ and $P_{n}^{\star}=1$.
Let $d_{i}:=n-i$ be the number of wild-type leaves. The three fitness totals are
$F_{i}^{\star}=ri+d_{i}+r,\qquad F_{i}^{\varnothing}=ri+d_{i}+1,\qquad G_{i}=ri+d_{i}.$ (79)
Following the analysis in [4], at a center-mutant state the process satisfies
$(1-C_{i})P_{i}^{\star}=A_{i}P_{i+1}^{\star}+B_{i}P_{i}^{\varnothing},$ (80)
where
$A_{i}$ $=(1-\delta)\frac{rd_{i}}{nF_{i}^{\star}}+\delta\frac{d_{i}}{N},$ (81)
$B_{i}$ $=(1-\delta)\frac{d_{i}}{F_{i}^{\star}}+\delta\frac{d_{i}}{NG_{i}},$ (82)
$C_{i}$ $=(1-\delta)\Bigl(\frac{ri}{F_{i}^{\star}}+\frac{ri}{nF_{i}^{\star}}\Bigr)+\delta\Bigl(\frac{i}{N}+\frac{ri}{NG_{i}}\Bigr).$ (83)
Solving: $P_{i}^{\star}=\alpha_{i}P_{i+1}^{\star}+\beta_{i}P_{i}^{\varnothing}$ with $\alpha_{i}=A_{i}/(1-C_{i})$ and $\beta_{i}=B_{i}/(1-C_{i})$.
At a center wild-type state,
$(1-b_{i})P_{i}^{\varnothing}=a_{i}P_{i}^{\star}+c_{i}P_{i-1}^{\varnothing},$ (84)
where
$a_{i}$ $=(1-\delta)\frac{ri}{F_{i}^{\varnothing}}+\delta\frac{ri}{NG_{i}},$ (85)
$c_{i}$ $=(1-\delta)\frac{i}{nF_{i}^{\varnothing}}+\delta\frac{i}{N},$ (86)
$b_{i}$ $=1-a_{i}-c_{i}.$ (87)
Solving: $P_{i}^{\varnothing}=p_{i}P_{i}^{\star}+q_{i}P_{i-1}^{\varnothing}$ with $p_{i}=a_{i}/(1-b_{i})$ and $q_{i}=c_{i}/(1-b_{i})$.
These two recurrences combine into the matrix form
$\begin{pmatrix}P_{i+1}^{\star}\\ P_{i}^{\varnothing}\end{pmatrix}=M_{i}\begin{pmatrix}P_{i}^{\star}\\ P_{i-1}^{\varnothing}\end{pmatrix},\quad M_{i}=\begin{pmatrix}\frac{1-\beta_{i}p_{i}}{\alpha_{i}}&-\frac{\beta_{i}q_{i}}{\alpha_{i}}\\ p_{i}&q_{i}\end{pmatrix}$. (88)
Setting $A^{(i)}=\prod_{j=1}^{i}M_{j}$, the boundary conditions give
$P_{1}^{\star}=1/A^{(n-1)}_{1,1}$, (89)
and then
$\mathsf{fp}_{r}^{\delta}(S_{N},\{c\})=P_{0}^{\star}=\alpha_{0}P_{1}^{\star},\qquad\mathsf{fp}_{r}^{\delta}(S_{N},\{\ell_{i}\})=P_{1}^{\varnothing}=p_{1}P_{1}^{\star}.$ (90)
When $r=1$, the coefficients simplify and the formulas in Theorem 3.6 are recovered (see Section 7).
12 Proofs of Theorems 3.12 and 3.13
Definition 12.1 (Almost regular).
An unweighted undirected graph is $\lambda$-almost regular if its maximum degree is at most $\lambda$ times its minimum degree.
Theorem 12.1.
For $\lambda\geq 1$, a $\lambda$-almost regular graph $G=(V,E)$ on $N$ vertices with $r\geq\lambda^{2}$, the mixed $\delta$-updating process has the following properties for all $\emptyset\subsetneq S\subsetneq V$:
  1. 1.
    $\mathsf{AT}_{r}^{\delta}(G,S)\leq\frac{r}{r-1}\cdot N^{5}$; and
  2. 2.
    $\mathsf{fp}_{r}^{\delta}(G,S)\geq|S|/N^{2}$.
Proof.
With $\phi(S)=|S|$, the expected one-step change satisfies
$\mathbb{E}\big[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S\big]=\sum_{u\in S}\sum_{v\notin S}\psi_{u,v}^{\delta}(S).$ (91)
We show $\psi_{u,v}^{\delta}(S)\geq\frac{r-1}{rN^{3}}$ for each crossing edge $(u,v)$ by splitting into three cases.
Case 1: $\delta=1/2$. By Theorem 9.1, $\psi_{u,v}^{\delta=1/2}(S)\geq\frac{r-1}{rN^{3}}$.
Case 2: $\delta<1/2$. Write $\psi_{u,v}^{\delta}(S)=(1-2\delta)\psi_{u,v}^{\mathrm{Bd}}(S)+2\delta\psi_{u,v}^{\delta=1/2}(S)$. We have
$\psi_{u,v}^{\mathrm{Bd}}(S)=\frac{1}{w(S)}\left(\frac{r}{\deg(u)}-\frac{1}{\deg(v)}\right)\geq\frac{1}{rN}\cdot\frac{r-\deg(u)/\deg(v)}{\deg(u)}\geq\frac{r-\lambda}{rN^{2}}.$ (92)
Since $r\geq\lambda^{2}$, we have $r-1\geq(\lambda-1)(\lambda+1)$, so $r-\lambda\geq(r-1)/N$ for $N\geq 2$, giving $\psi_{u,v}^{\mathrm{Bd}}(S)\geq\frac{r-1}{rN^{3}}$. Thus $\psi_{u,v}^{\delta}(S)\geq\frac{r-1}{rN^{3}}$.
Case 3: $\delta>1/2$. Use potential $\phi(S)=\sum_{u\in S}\deg(u)$ and write $\psi_{u,v}^{\delta}(S)=(2\delta-1)\psi_{u,v}^{\mathrm{dB}}(S)+2(1-\delta)\psi_{u,v}^{\delta=1/2}(S)$. For the dB term: if $u$ has a mutant neighbor or $v$ has a wild-type neighbor, then $\psi_{u,v}^{\textrm{dB}}(S)\geq\frac{r-1}{rN^{2}}\geq\frac{r-1}{rN^{3}}$; otherwise (both in boundary configurations) $\psi_{u,v}^{\textrm{dB}}(S)=0$ but $\psi_{u,v}^{\delta=1/2}(S)\geq\frac{r-1}{rN^{3}}$. In either case $\psi_{u,v}^{\delta}(S)\geq\frac{r-1}{rN^{3}}$.
Summing over crossing edges and applying Corollary 8.1 with $k_{1}=\sum_{v}\deg(v)\leq N^{2}$ and $k_{2}=\frac{r-1}{rN^{3}}$ proves item 1. For item 2, apply Theorem 8.2 with $k_{1}=|S|$ and $k_{2}=N^{2}$ (since $\phi(S)\leq N^{2}$ in both regimes), giving $\mathsf{fp}_{r}^{\delta}(G,S)\geq|S|/N^{2}$. ∎
Proof of Theorem 3.12.
Set $\lambda=D/d$ in Theorem 12.1. Since $r>(D/d)^{2}=\lambda^{2}$ and the graph is $\lambda$-almost regular, item 2 gives $\mathsf{fp}_{r}^{\delta}(G,\{u\})\geq 1/N^{2}$. ∎
Proof of Theorem 3.13.
Set $\lambda=D/d$ in Theorem 12.1. Since $r>\lambda^{2}$ and $\mathsf{fp}_{r}^{\delta}(G,S)\geq|S|/N^{2}\geq 1/N^{2}$ by Theorem 3.12, we get $\mathsf{FT}_{r}^{\delta}(G,\{u\})\leq\mathsf{AT}_{r}^{\delta}(G,\{u\})/\mathsf{fp}_{r}^{\delta}(G,\{u\})\leq\frac{r}{r-1}\cdot N^{5}\cdot N^{2}=\frac{r}{r-1}\cdot N^{7}$. ∎
13 FPRAS Amplification
Theorem 13.1 (FPRAS powering lemma).
If there is an FPRAS with confidence 1c1-c for some constant $c>0$, then there is an FPRAS with confidence $1-\alpha$ for $\alpha<c$.
Proof.
The idea is to run the FPRAS $b_{c}\cdot\lceil\log(1/\alpha)\rceil+1$ times for some particular constant $b_{c}>0$. Then return the median of the runs. See Lemma 6.1 in [56] for the proof and how to choose $b_{c}$. ∎
14 FPRAS for $\delta=1/2$
Theorem 14.1 (FPRAS for $\delta=1/2$).
There exists a polynomial time (with respect to $N$, $1/\varepsilon$, and $\log(1/\alpha)$) randomized algorithm $\mathcal{A}$ such that on input $r>1$, $\varepsilon>0$, $\alpha\in(0,1]$, an undirected unweighted graph $G=(V,E)$ with $N$ vertices, and $S_{0}\subseteq V$, outputs with probability at least $1-\alpha$ an estimate to the fixation probability within a relative error of $\varepsilon$. In other words,
$\mathbb{P}\left\{\big|\mathcal{A}(G,S_{0},r)-\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\big|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\right\}<\alpha.$ (93)
The probability is over sampling the randomness for $\mathcal{A}$.
Proof.
The algorithm that follows uses a similar technique to that of Theorem 3.14. Let $s=\lceil 2\ln(2/\alpha_{1})/(\rho\varepsilon)^{2}\rceil$ where $\rho=|S_{0}|/N$ for $\alpha_{1}=1/8$. Let $T=\lceil\alpha_{2}^{-1}sN^{4}r/(r-1)\rceil$ for $\alpha_{2}=1/8$.
Consider $s$ independent copies of the mixed $1/2$-updating process that will run all the way to absorption. Let $\mathcal{E}_{2}$ be the event that at least one copy will not absorb within $T$ steps. First notice that by a union bound, Markov’s inequality, and Theorem 9.1, we have $\mathbb{P}\{\mathcal{E}_{2}\}<s\cdot\mathsf{AT}_{r}^{\delta=1/2}(G,S_{0})/T<\alpha_{2}$. If $\mathcal{E}_{2}$ occurs, we output 0 as our estimate for $\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})$. Otherwise all copies will absorb. For each $i\in\{1,\dots,s\}$, let $Z_{i}$ be the indicator that the $i$-th copy will fixate, and set $\hat{f}:=\frac{1}{s}\sum_{i=1}^{s}Z_{i}$. Then $Z_{1},\dots,Z_{s}$ are i.i.d. Bernoulli random variables with $\mathbb{E}[Z_{i}]=\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})$. Moreover, on the event $\neg\mathcal{E}_{2}$, all $s$ copies will absorb by time $T$; so in that case the algorithm observes all absorption outcomes and outputs $\hat{f}$. Hence the event FAIL that the algorithm’s output deviates from $\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})$ by more than a multiplicative factor $\varepsilon$ satisfies
$\textsf{FAIL}\subseteq\left(\mathcal{E}_{2}\cup\Big\{|\hat{f}-\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\Big\}\right).$ (94)
By a Chernoff bound and Corollary 9.1, we obtain
$\mathbb{P}\Big\{|\hat{f}-\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta=1/2}(G,S_{0})\Big\}<2\exp(-s\cdot(\rho\varepsilon)^{2}/2)<\alpha_{1}.$ (95)
Therefore by a union bound, the probability that our algorithm errs is at most
$\mathbb{P}\{\mathsf{FAIL}\}<\mathbb{P}\{\mathcal{E}_{2}\}+\alpha_{1}<\alpha_{1}+\alpha_{2}=1/4.$ (96)
We can check if $\mathcal{E}_{2}$ occurs by simulating each process for $T$ time steps and checking if any process has not absorbed. The running time of this algorithm is polynomial in $s$ and $T$, which are both polynomial in $N$ and $1/\varepsilon$. Finally, we use Theorem 13.1 which contributes a multiplicative $O(\log(1/\alpha))$ factor to our algorithm time. Thus the running time of the complete algorithm is polynomial in $N$, $1/\varepsilon$, and $\log(1/\alpha)$. ∎
15 Proof of Theorem 3.14
Proof of Theorem 3.14.
The algorithm simulates the mixed $\delta$-updating process many times and records how many trajectories fixated. Let $\rho=|S_{0}|/N^{2}$ (a lower bound on the fixation probability from Theorem 3.12). Let $s=\lceil 2\ln(2/\alpha_{1})/(\rho\varepsilon)^{2}\rceil$ where $\alpha_{1}=1/8$. Let $T=\lceil\alpha_{2}^{-1}sN^{5}r/(r-1)\rceil$ where $\alpha_{2}=1/8$.
Consider $s$ independent copies of the mixed $\delta$-updating process that will run all the way to absorption. Let $\mathcal{E}_{2}$ be the event that at least one copy will not absorb within $T$ steps. By a union bound, Markov’s inequality, and Theorem 12.1,
$\mathbb{P}\{\mathcal{E}_{2}\}<s\cdot\mathsf{AT}_{r}^{\delta}(G,S_{0})/T<\alpha_{2}.$ (97)
If $\mathcal{E}_{2}$ occurs, we output 0 as our estimate for $\mathsf{fp}_{r}^{\delta}(G,S_{0})$. Otherwise all copies absorb. For each $i\in\{1,\dots,s\}$, let $Z_{i}$ be the indicator that the $i$-th copy fixates, and set $\hat{f}:=\frac{1}{s}\sum_{i=1}^{s}Z_{i}$. Then $Z_{1},\dots,Z_{s}$ are i.i.d. Bernoulli random variables with $\mathbb{E}[Z_{i}]=\mathsf{fp}_{r}^{\delta}(G,S_{0})$. The event FAIL that the output deviates from $\mathsf{fp}_{r}^{\delta}(G,S_{0})$ by more than $\varepsilon$ satisfies
$\textsf{FAIL}\subseteq\left(\mathcal{E}_{2}\cup\Big\{|\hat{f}-\mathsf{fp}_{r}^{\delta}(G,S_{0})|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta}(G,S_{0})\Big\}\right).$ (98)
By a Chernoff bound and Theorem 3.12,
$\mathbb{P}\Big\{|\hat{f}-\mathsf{fp}_{r}^{\delta}(G,S_{0})|>\varepsilon\cdot\mathsf{fp}_{r}^{\delta}(G,S_{0})\Big\}<2\exp(-s\cdot(\rho\varepsilon)^{2}/2)<\alpha_{1}.$ (99)
Therefore
$\mathbb{P}\{\mathsf{FAIL}\}<\mathbb{P}\{\mathcal{E}_{2}\}+\alpha_{1}<\alpha_{1}+\alpha_{2}=1/4.$ (100)
The running time is polynomial in $s$ and $T$, both polynomial in $N$ and $1/\varepsilon$. Applying Theorem 13.1 contributes a multiplicative $O(\log(1/\alpha))$ factor, giving total runtime polynomial in $N$, $1/\varepsilon$, and $\log(1/\alpha)$. ∎